# Learning normalized image densities via dual score matching

### v1

@@ Abstract
Learning probability models from data is at the heart of many machine learning endeavors, but is notoriously difficult due to the curse of dimensionality. We introduce a new framework for learning normalized energy (log probability) models that is inspired from diffusion generative models, which rely on networks optimized to estimate the score. We modify a score network architecture to compute an energy while preserving its inductive biases. The gradient of this energy network with respect to its input image is the score of the learned density, which can be optimized using a denoising objective. Importantly, the gradient with respect to the noise level provides an additional score that can be optimized with a novel secondary objective, ensuring consistent and normalized energies across noise levels. We train an energy network with this dual score matching objective on the ImageNet64 dataset, and obtain a cross-entropy (negative log likelihood) value comparable to the state of the art. We further validate our approach by showing that our energy model strongly generalizes: estimated log probabilities are nearly independent of the specific images in the training set. Finally, we demonstrate that both image probability and dimensionality of local neighborhoods vary significantly with image content, in contrast with traditional assumptions such as concentration of measure or support on a low-dimensional manifold.
@@ 1 Introduction
Many problems in image processing and computer vision rely, explicitly or implicitly, on prior probability models. However, learning such models by maximizing the likelihood of a set of training images is difficult. The dimensionality of the space (i.e., number of image pixels) is large, and worst-case data requirements for estimation grow exponentially (the “curse of dimensionality”). The machine learning community has developed a variety of methods to train a parametric network to estimate $\log p(x)$, known as an “energy model” (Hinton et al., 1986; LeCun et al., 2006), relying on the inductive biases of the network to alleviate the data requirements. For all but the simplest of models, this approach is frustrated by the intractability of estimating the normalization constant (Hinton, 2002; LeCun and Huang, 2005; Yedidia et al., 2005; Gutmann and Hyvärinen, 2010; Dinh et al., 2014; Rezende and Mohamed, 2015; Dinh et al., 2017; Song and Kingma, 2021).
@@
A clever means of escaping this conundrum is to estimate the gradient of the energy with respect to the image (known as the “score”), which eliminates the normalization constant, and can be learned from data with a “score-matching” objective (Hyvärinen and Dayan, 2005). The recent development of “diffusion” generative models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Kadkhodaie and Simoncelli, 2020) builds on this concept, by estimating a family of score functions for images corrupted by Gaussian white noise at different amplitudes. The scores may then be used to sample from the corresponding estimated density, using an iterative reverse diffusion procedure. These methods have enabled dramatic improvements in both the quality and diversity of generated image samples, but the learned density is implicit. An explicit and normalized energy model (and density) can be obtained through integration of the divergence of the score vector field along a trajectory, at tractable but large computational cost.
@@
Here, we leverage the power of diffusion models to develop a robust and efficient framework for directly learning a normalized energy model from image data. We approximate the energy with a deep neural network that takes as input both a noisy image and the corresponding noise variance, and derive two separate objectives by differentiating with respect to these two inputs. The first is a denoising objective, as used in diffusion models. The second is novel and ensures consistency of the energy estimates across noise levels, which we show to be critical for obtaining accurate and normalized energies. We optimize the sum of the two, in a procedure that we refer to as “dual score matching”. We also propose a novel architecture for energy estimation by computing the inner product of the output of a score network with its input image. This preserves the inductive biases of the base score network, leading to equal or superior denoising performance (and thus sample quality).
@@
We train our energy model on ImageNet64 (Russakovsky et al., 2015; Chrabaszcz et al., 2017), and show that the estimated energies lead to negative log likelihood (or cross-entropy) values comparable to the state-of-the-art. We further demonstrate that the energy model strongly generalizes in the sense of Kadkhodaie et al. (2024): two separate models trained on non-overlapping subsets of the training data assign the same probabilities to a given image. This convergence is observed at moderate training set sizes, and is far faster than the worst case prediction of the curse of dimensionality. We find that the distribution of log probabilities over ImageNet images covers a broad range, with densely textured images at the lower end, and sparse images at the upper end. The probability is stable with respect to image luminance, but decreases with dynamic range. Finally, we highlight two geometrical properties of the learned image distribution. The first is an extremely tight inverse relationship between volume and density that leads to an absence of concentration of energy values. They furthermore follow a Gumbel distribution, indicating a surprising statistical regularity. The second is that the local dimensionality of the energy landscape in the neighborhood of an image varies greatly depending on image content and neighborhood size. We find both images with full-dimensional neighborhoods of non-negligible size and images with lower-dimensional neighborhoods even at sub-quantization scales. These results challenge several traditional presuppositions regarding high-dimensional distributions, such as the concentration of measure phenomenon (Vershynin, 2018; Wainwright, 2019) or the manifold hypothesis (Tenenbaum et al., 2000; Bengio et al., 2013). We release code to reproduce all experiments and pre-trained models at https://github.com/FlorentinGuth/DualScoreMatching.
@@ 2 Learning normalized energy models with dual score matching
Estimating a high-dimensional probability density from samples faces two significant challenges as a result of the curse of dimensionality. The first is statistical: reasonably-sized datasets do not contain enough information about the unknown distribution. One then needs powerful inductive biases, typically in the form of a network architecture, to hope to recover the data distribution. The second challenge is computational: the traditional objective is to maximize the likelihood of the model over the data, which is intractable due to the need to compute the normalization constant. Nevertheless, recently developed generative models implicitly solve this problem. Our approach draws inspiration from diffusion models (Section 2.1), and derives a novel objective (Section 2.2) and architecture (Section 2.3) to learn normalized log probabilities (energies) from data, addressing both challenges. It is validated numerically in Section 2.4.
@@ 2.1 Motivation
Figure 1: Comparison of single and dual score matching on recovering the energy of a scale mixture of two Gaussians in $d=1000$ dimensions. Left: Radial slices of the log probability. The single score matching estimate (green dashed curve) fails to recover the true energy (blue solid curve), even after global normalization (green dotted curve), while dual score matching (red dashed curve) succeeds. Middle: Radial components of the scores. Single score matching learns an accurate score over the support of the data (blue bar plot) but not outside of it. Right: Energy landscape across space and time (noise level) for a mixture of two Gaussians in one dimension. The direct path between the modes at $t=0$ crosses a large energy barrier (green curve), which is alleviated on a path that is not time-restricted (red curve).
@@
Traditionally, energy models are defined in terms of a parametric function $U_{\theta}(x)$ that approximates the unnormalized log density over $x\in\mathbb{R}^{d}$: $p_{\theta}(x)=\frac{1}{Z_{\theta}}\mathrm{e}^{-U_{\theta}(x)}$, with a normalizing constant $Z_{\theta}=\int\mathrm{e}^{-U_{\theta}(x)}\mathrm{d}x$. The parameters $\theta$ are estimated by minimizing the expected negative log likelihood (NLL), $\mathbb{E}_{x}\mathopen{}\mathclose{{}\left[-\log p_{\theta}(x)}\right]=\mathbb{E}_{x}\mathopen{}\mathclose{{}\left[U_{\theta}(x)}\right]+\log Z_{\theta}$, which is equivalent to minimizing the KL divergence between $p_{\theta}(x)$ and the data distribution. The normalizing constant $Z_{\theta}$ plays a critical role in learning, representing the total energy over $\mathbb{R}^{d}$ which trades off against the energy of the data. But direct estimation (i.e., computing the integral) is typically intractable.
@@
The normalization constant can be eliminated from the NLL by differentiating w.r.t. $x$, yielding a quantity known as the score: $-\nabla_{x}\log p_{\theta}(x)=\nabla_{x}U_{\theta}(x)$. As a result, a score model can be efficiently fitted to data via “score matching” (Hyvärinen and Dayan, 2005), which minimizes the Fisher (rather than KL) divergence between the model and the data. In the case of data corrupted by Gaussian white noise, this amounts to solving a denoising problem (Vincent, 2011; Raphan and Simoncelli, 2011). But this computational advantage comes at a statistical cost: a good approximation of the score does not always lead to a good model of the energy (Koehler et al., 2023), as we now illustrate. Consider the case of an equal mixture of two multivariate Gaussian distributions with zero mean and different variances $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$. In high dimensions, this distribution concentrates near the two spheres of radius $\sigma_{1}\sqrt{d}$ and $\sigma_{2}\sqrt{d}$, leading to data scarcity in the rest of the space. We show in Figure 1 energies estimated from samples by optimizing their gradient via (single) score matching. This fails to recover the true energy, even after adjusting a normalization constant. Indeed, estimating the energy difference between the two Gaussians requires a good estimation of the score along an integration path between them. If there is no data in-between modes to constrain the score, due to an energy barrier or concentration phenomena, this leads to inconsistent energy values.
@@
Diffusion models expand on score matching by learning scores of data corrupted by white Gaussian noise for a range of different noise amplitudes, $\nabla_{y}U(y,t)$, with $y=x+\mathcal{N}(0,t\mathrm{Id})$. The evolution of the density $p(y|t)$ as $t$ increases is a diffusion process, and thus the corresponding energies increase in smoothness with $t$. This multiscale family of scores can be used to draw high-quality samples from $p(x)$ using a reverse diffusion algorithm, which follows a trajectory of partial denoising steps (Song and Ermon, 2019; Song et al., 2021a; Ho et al., 2020; Kadkhodaie and Simoncelli, 2020). In fact, the diffusion scores implicitly capture a density model of the data (Song et al., 2021b; Kingma and Welling, 2013): the relative energy levels between modes are encoded in the score at the time $t$ when they merge (Raya and Ambrogioni, 2023; Biroli et al., 2024). This density model can be evaluated in various ways, which all involve a tractable but costly integration of score divergences (or denoising errors) along time (Song et al., 2021a; Kong et al., 2023; Skreta et al., 2024; Karczewski et al., 2025). Intuitively, diffusion provides a high-probability path between modes through space and time, as visualized in Figure 1. Energy differences along time can then be retrieved by integrating the “time score” $\partial_{t}U_{\theta}(y,t)$ (Choi et al., 2022), which is implicitly determined from the diffusion equation (Song et al., 2021a; Lai et al., 2023). See Section A.2 for a more in-depth exposition. Here, we propose to learn the “space-time” energy landscape directly by matching both space and time scores.
@@ 2.2 Dual score matching: Objective function
We want to learn a time-dependent energy function, $U_{\theta}(y,t)$, to approximate the NLL of the noisy image distribution at all noise levels $t$:
@@
Differentiating this NLL with respect to $y$ gives the “space score”, which can be expressed using the Miyasawa-Tweedie identity (Robbins, 1956; Miyasawa, 1961) as
@@
It leads to the denoising score matching objective (Vincent, 2011; Raphan and Simoncelli, 2011) used to train diffusion models:
@@
We also can differentiate the energy with respect to $t$, producing a “time score”, for which we can derive a similar score-matching identity (see Section B.1):
@@
From this identity, we define an analogous “time score matching” objective:
@@
Matching both scores enforces that the learned energy $U_{\theta}(y,t)$ is a solution of the diffusion equation, which is critical for combining information across noise levels into accurate energy values at $t=0$.
@@
The two objectives in eqs. 3 and 5 are defined for a fixed noise level $t$. We can form an overall objective by integrating over $t$, with an appropriate weighting. For the denoising score matching objective, a natural choice is the so-called maximum-likelihood weighting (Song et al., 2021b; Kingma et al., 2021), which provides a bound on the KL divergence with the data distribution:
@@
Appreciably, the resulting term in the expected value is unitless: it is invariant to a simultaneous rescaling of the data and noise level. We thus choose to weight the time score matching objective in eq. 5 by $t^{2}$, so that it is also unitless, and to evaluate over the same distribution $p(t)$. Finally, after normalization by the dimensionality $d$, we simply add both objectives (we found no significant improvement from tuning a tradeoff hyperparameter). Our dual score matching objective is then
@@
The learned energy $U_{\theta}(y,t)$ is only estimated up to an additive constant, since the two objective functions only constrain its partial derivatives. An important aspect of our framework is that it enables estimation of this normalizing constant. Indeed, the time score objective ensures that this normalizing constant does not depend on time, as mass is conserved through the diffusion. This allows normalizing the model at the largest noise level $t_{\max}$, where the true distribution is almost Gaussian: $p(y|t_{\max})\approx\mathcal{N}(0,t_{\max}\mathrm{Id})$. We thus set the average energy value at $t=t_{\max}$ to the entropy of this Gaussian distribution:
@@
Figure 1 verifies that dual score matching indeed provides both consistent and normalized energy values (in particular, for $t=0$), in contrast to single score matching. We note however that eq. 9 is an approximate normalization procedure, which we discuss at more length in Section A.3.

### v3

@@ Abstract
Learning probability models from data is at the heart of many machine learning endeavors, but is notoriously difficult due to the curse of dimensionality. We introduce a new framework for learning normalized energy (log probability) models that is inspired by diffusion generative models, which rely on networks optimized to estimate the score. We modify a score network architecture to compute an energy while preserving its inductive biases. The gradient of this energy network with respect to its input image is the score of the learned density, which can be optimized using a denoising objective. Importantly, the gradient with respect to the noise level provides an additional score that can be optimized with a novel secondary objective, ensuring consistent and normalized energies across noise levels. We train an energy network with this dual score matching objective on the ImageNet64 dataset, and obtain a cross-entropy (negative log likelihood) value comparable to the state of the art. We further validate our approach by showing that our energy model strongly generalizes: log probabilities estimated with two networks trained on non-overlapping data subsets are nearly identical. Finally, we demonstrate that both image probability and dimensionality of local neighborhoods vary substantially depending on image content, in contrast with conventional assumptions such as concentration of measure or support on a low-dimensional manifold.
@@ 1 Introduction
Many problems in image processing and computer vision rely, explicitly or implicitly, on prior probability models. However, learning such models by maximizing the likelihood of a set of training images is difficult. The dimensionality of the space (i.e., the number of image pixels) is large, and worst-case data requirements for estimation grow exponentially (the “curse of dimensionality”). The machine learning community has developed a variety of methods to train a parametric network to estimate $\log p(x)$, known as an “energy model” (Hinton et al., 1986; LeCun et al., 2006), relying on the inductive biases of the network to alleviate the data requirements. For all but the simplest of models, this approach is frustrated by the intractability of estimating the normalization constant (Hinton, 2002; LeCun and Huang, 2005; Yedidia et al., 2005; Gutmann and Hyvärinen, 2010; Dinh et al., 2014; Rezende and Mohamed, 2015; Dinh et al., 2017; Song and Kingma, 2021).
@@
A clever means of escaping this conundrum is to estimate the gradient of the energy with respect to the image (known as the “score”), which eliminates the normalization constant, and can be learned from data with a “score-matching” objective (Hyvärinen and Dayan, 2005). The recent development of “diffusion” generative models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Kadkhodaie and Simoncelli, 2020) builds on this concept, by estimating a family of score functions for images corrupted by Gaussian white noise at different amplitudes. These scores may then be used to draw samples from the corresponding estimated density, using an iterative reverse diffusion procedure. These methods have enabled dramatic improvements in both the quality and diversity of generated image samples, but the learned density is implicit. An explicit and normalized energy model (and density) can be obtained through integration of the divergence of the score vector field along a trajectory (Song et al., 2021a), at tractable but substantial computational cost.
@@
Here, we leverage the power of diffusion models to develop a robust and efficient framework for directly learning a normalized energy model from image data. We approximate the energy with a deep network that takes as input both a noisy image and the corresponding noise variance, and derive two separate objectives. The first, obtained by differentiating with respect to the noisy image, is a denoising objective (as used in diffusion models). The second, obtained by differentiating with respect to the noise variance, ensures consistency of the energy estimates across noise levels, which we show to be critical for obtaining accurate and normalized energies. We optimize the sum of the two, in a procedure that we refer to as “dual score matching”. We also propose a novel architecture for energy, computed as the inner product of the input and output images of a score network. This preserves the inductive biases of the base score network, leading to equal or superior denoising performance (and thus, sample quality).
@@
We train our energy model on ImageNet64 (Russakovsky et al., 2015; Chrabaszcz et al., 2017), and show that the estimated energies lead to negative log likelihood (or cross-entropy) values comparable to the state of the art. We further demonstrate that the energy model strongly generalizes in the sense of Kadkhodaie et al. (2024): two separate models trained on non-overlapping subsets of the training data assign essentially the same probabilities to each image. This convergence is observed for training set sizes far smaller than the worst case prediction of the curse of dimensionality. We find that the distribution of log probabilities over ImageNet images covers a broad range, with densely textured images at the lower end, and sparse images at the upper end. The probability is relatively insensitive to changes in image luminance, but decreases with dynamic range. Finally, we highlight two geometrical properties of the learned image distribution. The first is an extremely tight inverse relationship between volume and density that leads to an absence of concentration of energy values. They furthermore follow a Gumbel distribution, revealing an unexpected form of statistical regularity. The second is that the dimensionality of the energy landscape in the neighborhood of an image varies greatly depending on image content and neighborhood size. We observe images with full-dimensional neighborhoods of non-negligible size and images with lower-dimensional neighborhoods even at sub-quantization scales. These results challenge traditional presuppositions regarding high-dimensional distributions, such as the concentration of measure phenomenon (Vershynin, 2018; Wainwright, 2019) and the manifold hypothesis (Tenenbaum et al., 2000; Bengio et al., 2013). Pre-trained models and software for running all experiments are available at https://github.com/FlorentinGuth/DualScoreMatching.
@@ 2 Learning normalized energy models with dual score matching
Estimating a high-dimensional probability density from samples faces two significant challenges as a result of the curse of dimensionality. The first is statistical: realistically-sized datasets do not contain enough information to learn an unknown high-dimensional distribution. One thus needs powerful inductive biases, typically in the form of a parametric network architecture, to hope to recover the data distribution. The second challenge is computational: the traditional objective aims to maximize the likelihood of the model over the data, which is intractable due to the need to compute the normalization constant. Nevertheless, recently developed generative models seem to succeed despite these problems. Drawing inspiration from diffusion models (Section˜2.1), we develop a more direct solution by deriving a novel objective (Section˜2.2) and architecture (Section˜2.3) for learning normalized log probabilities (energies) from data. We validate our approach in Section˜2.4.
@@ 2.1 Motivation
Figure 1: Comparison of single and dual score matching on recovering the energy of a scale mixture of two Gaussians in $d=1000$ dimensions. Experimental details are provided in Section˜C.3. Left: Radial slices of the log probability. The single score matching estimate (green dashed curve) fails to recover the true energy (blue solid curve), even after global normalization (green dotted curve), while dual score matching (red dashed curve) succeeds. Middle: Radial components of the scores. Single score matching learns an accurate score over the support of the data (blue bar plot) but not outside of it. Right: Energy landscape across space and time (noise level) for a mixture of two Gaussians in one dimension. The direct path between the modes at $t=0$ crosses a large energy barrier (green curve), which is alleviated on a path that is not restricted to $t=0$ (red curve).
@@
Traditionally, energy models are defined in terms of a parametric function $U_{\theta}(x)$ that approximates the unnormalized log density over $x\in\mathbb{R}^{d}$: $p_{\theta}(x)=\frac{1}{Z_{\theta}}\mathrm{e}^{-U_{\theta}(x)}$, with a normalizing constant $Z_{\theta}=\int\mathrm{e}^{-U_{\theta}(x)}\mathrm{d}x$. The parameters $\theta$ are estimated by minimizing the expected negative log likelihood (NLL), $\mathbb{E}_{x}\mathopen{}\mathclose{{\left[-\log p_{\theta}(x)}}\right]=\mathbb{E}_{x}\mathopen{}\mathclose{{\left[U_{\theta}(x)}}\right]+\log Z_{\theta}$, which is equivalent to minimizing the KL divergence between $p_{\theta}(x)$ and the data distribution. The normalizing constant $Z_{\theta}$ plays a critical role in learning, representing the total energy over $\mathbb{R}^{d}$ which trades off against the energy of the data. Unfortunately, direct estimation (i.e., computing the integral) is typically intractable.
@@
The normalization constant can be eliminated from the NLL by differentiating w.r.t. $x$, yielding a quantity known as the (negative) score: $\nabla_{x}\log p_{\theta}(x)=\nabla_{x}U_{\theta}(x)$. As a result, a score model can be efficiently fitted to data via “score matching” (Hyvärinen and Dayan, 2005), which minimizes the Fisher (as opposed to KL) divergence between the model and the data. For data corrupted by Gaussian white noise, this amounts to solving a denoising problem (Vincent, 2011; Raphan and Simoncelli, 2011). But this computational advantage comes at a statistical cost: a good approximation of the score does not always lead to a good model of the energy (Koehler et al., 2023), as we now illustrate. Consider the case of an equal mixture of two multivariate Gaussian distributions with zero mean and variances $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$. In high dimensions, this distribution concentrates near the two hyper-spheres with radii $\sigma_{1}\sqrt{d}$ and $\sigma_{2}\sqrt{d}$, leading to data scarcity in the rest of the space. We show in Figure˜1 energies estimated from samples by optimizing their gradient via (single) score matching. This fails to recover the true energy, even after adjusting a normalization constant. Indeed, evaluating the energy difference between the two Gaussians requires good estimation of the score along an integration path between them. If there is no data in-between modes to constrain the score, due to an energy barrier or concentration phenomena, this leads to inconsistent energy values across modes. In other words, single score matching estimates energy values up to a mode-dependent additive constant.
@@
Diffusion models expand on score matching by learning scores of data corrupted by white Gaussian noise for a range of different noise amplitudes, $\nabla_{y}U(y,t)$, with $y=x+\mathcal{N}(0,t\mathrm{Id})$. The evolution of the density $p(y|t)$ as “time” (noise variance) $t$ increases is a diffusion process, and thus the corresponding energies increase in smoothness with $t$. This multiscale family of scores can be used to draw high-quality samples from $p(x)$ using a reverse diffusion algorithm, which follows a trajectory of partial denoising steps (Song and Ermon, 2019; Song et al., 2021a; Ho et al., 2020; Kadkhodaie and Simoncelli, 2020). In fact, the diffusion scores implicitly capture a density model of the data (Song et al., 2021b; Kingma et al., 2021): the relative energy levels between modes, to which score matching at $t=0$ is blind (Zhang et al., 2022), are encoded in the score at the time $t$ when they merge (Raya and Ambrogioni, 2023; Biroli et al., 2024). This implicit density model can be evaluated in various ways, which all involve a tractable but costly integration of score divergences (or denoising errors) as a function of time (Song et al., 2021a; Kong et al., 2023; Skreta et al., 2024; Karczewski et al., 2025). See Section˜A.2 for a more in-depth exposition.
Intuitively, diffusion models deal with multimodal distributions by providing a high-probability path between modes through space and time, as visualized in the right panel of Figure˜1. In other words, the joint distribution $p(y,t)$ qualitatively has a connected support. This suggests that we may learn the joint “space-time” energy $U(y,t)$ by score matching on $(y,t)$. Specifically, the “space score” $\nabla_{y}U_{\theta}(y,t)$ can be learned using a denoising objective, and the “time score” $\partial_{t}U_{\theta}(y,t)$ (Choi et al., 2022) can be learned using an analogous score matching objective. Matching both space and time scores correctly constrains the energy levels across modes: the energy difference between two modes can be explicitly recovered by integrating the energy derivative along the red path in Figure˜1 (although we will not need to do so explicitly).
@@ 2.2 Dual score matching: Objective function
We aim to learn a time-dependent energy function, $U_{\theta}(y,t)$, to approximate the NLL of the noisy image distribution at all noise levels $t$:
@@
Differentiating this NLL with respect to $y$ gives the (negative) “space score”, which can be expressed using the Miyasawa-Tweedie identity (Robbins, 1956; Miyasawa, 1961) as
@@
This leads to the denoising score matching objective (Vincent, 2011; Raphan and Simoncelli, 2011; Saremi et al., 2018) used to train diffusion models:
@@
Differentiating the energy with respect to $t$ yields a (negative) “time score” (Choi et al., 2022), for which we can derive a similar identity (see Section˜B.1):
@@
This leads to an analogous “time score matching” objective:
@@
The two objectives in eqs.˜3 and 5 are defined for a given noise level $t$. We form an overall objective by integrating these over $t$, with an appropriate weighting. For the denoising score matching objective, a natural choice is the so-called maximum-likelihood weighting (Song et al., 2021b; Kingma et al., 2021), which provides a bound on the KL divergence with the data distribution:
@@
Note that the resulting term in the expected value is unit-less: it is invariant to simultaneous rescaling of the data and noise level. We thus choose to weight the time score matching objective in eq.˜5 by $t^{2}$, so that it is also unit-less, and to evaluate it over the same distribution $p(t)$. Finally, after appropriate normalization by the dimensionality $d$ that ensures that the two objectives have comparable orders of magnitude, we simply add them (we found no significant improvement from tuning a tradeoff hyperparameter). In summary, our dual score matching objective is:
@@
Since we match only the partial derivatives of $U_{\theta}(y,t)$ to those of the true energy $U(y,t)$, we only recover the energy up to a global constant: at the end of training, $U_{\theta}(y,t)\approx U(y,t)+\mathrm{const}$. Note again that this crucially relies on $p(y,t)$ having a connected support. An important aspect of our framework is that it enables estimation of this constant, which determines the normalization of $\mathrm{e}^{-U_{\theta}(y,t)}$. Indeed, the time score objective ensures that this normalizing constant does not depend on time: mass is conserved through the diffusion. Since the true distribution is approximately Gaussian at large noise levels, $p(y|t_{\max})\approx\mathcal{N}(0,t_{\max}\mathrm{Id})$, we can set this constant to the entropy of this Gaussian distribution:
@@
Figure˜1 provides a high-dimensional numerical example verifying that dual score matching indeed provides an accurate estimate of normalized energy values (in particular, for $t=0$), in contrast to single score matching.
Related approaches.
Similar combinations of space and time scores were considered by Choi et al. (2022) (who referred to it as the “pathwise” method) and Kobler and Pock (2023). These time score objectives, however, relied on a second derivative in time instead of a regression objective. Our time score objective is also a special case of the conditional time score matching objective developed concurrently in Yu et al. (2025). Finally, Yadin et al. (2024) train an energy model with an objective combining score matching with a classification cross-entropy loss for estimating a discretized version of the noise level $t$.